OpenStudy (anonymous):

Divide: (2x3 - x2 + 3x -1) ÷ (x + 2)

OpenStudy (anonymous):

\[\frac{2 x^3-x^2+3 x-1}{x+2}=2 x^2-5 x+13-\frac{27}{x+2} \]